OpenStudy (anonymous):

Solve the slope-intercept formula, y =mx+b for m A.m=y+b/x B.m=x(y-b) C.m=yx-b D.m=y-b/x

OpenStudy (alexandervonhumboldt2):

y=mx+b y-b=mx m=(y-b)/x

OpenStudy (alexandervonhumboldt2):

m=(y-b)/x

OpenStudy (anonymous):

there is no answer for that

OpenStudy (alexandervonhumboldt2):

m=y-b/x

OpenStudy (alexandervonhumboldt2):

remove ( )

OpenStudy (anonymous):

?

OpenStudy (alexandervonhumboldt2):

revocve bractets D

OpenStudy (anonymous):

oh srry im not very smart

OpenStudy (anonymous):

thx